OpenStudy (help!!!!):

To win an election, Jim need 3/4 of votes. If after 2/3 of votes have been counted Edith has 5/6 of what he needs, what part of the remaining votes does she still need?

OpenStudy (dan815):

he currently has 5/6 of 3/4

OpenStudy (dan815):

15/24, he needs 3/4 votes to win or 18/24 votes to win

OpenStudy (help!!!!):

The answer choices are 1/8, 3/8, 5/8, 5/24

OpenStudy (dan815):

xD

OpenStudy (dan815):

okay so think about it

OpenStudy (dan815):

what fraction of 1/3 does he need to complete 3/4

OpenStudy (dan815):

15/24 + f*(1/3) = 3/4

OpenStudy (help!!!!):

3/8 :)

MsBrains (ms-brains):

f = 3/8

OpenStudy (dan815):

yes

OpenStudy (dan815):

does it make sense?

OpenStudy (help!!!!):

yup! Thanks

OpenStudy (dan815):

ok sure!

OpenStudy (help!!!!):

wait actually how did you get 15/24? @dan815

OpenStudy (dan815):

5/6 of 3/4 5/6 * 3/4

OpenStudy (help!!!!):

hmm okey just multiply the # you need and the # you have

OpenStudy (help!!!!):

plus the # left over * x. Yeah I got it!

OpenStudy (campbell_st):

apart from having the names wrong... if Jim needs 3/4 does that mean Edith wins with any votes... ?

OpenStudy (campbell_st):

so if Edith has 5/6 pf 2/3 then Jim can't win...

OpenStudy (help!!!!):

ops typo is the same person. "Jim". I just switched out the names

OpenStudy (campbell_st):

ok... so if jim has 5/6 of 2/3 then he has 10/18 or 20/36 he needs 27/36 in total so he needs 7/36 from the remaining 1/3 of the votes... so I think you'll find Jim needs to win 7/12 of the remaining votes

OpenStudy (campbell_st):

but, that's just my opinion...

OpenStudy (dan815):

oh lol i didnt even notice edith is a new name there

OpenStudy (dan815):

i just looked at the numbers and assumed the question hehe

OpenStudy (campbell_st):

do you think you got the number right..?

OpenStudy (dan815):

ya it sshould be fine so another way to think about it is in term marks say u need 75% or 3/4 to pass a course currently you got 5/6 of the 75 marks so 62.5 2/3 is 66.66 62.5/66.66 is like 93% u currently have a 93% avg in the course but anyay the point is u need another 12.5=75/62.5 12.5 from 33.33 12.5 is what fraction of 33.33

OpenStudy (dan815):

which is 3/8

OpenStudy (campbell_st):

no... you have 5/6 of votes when 2/3 of the total votes counted,,, they haven't counted 3/4 of the votes...

OpenStudy (dan815):

it's 5/6 of the total votes he needs after 2/3 its 5/6 * 3/4 not 5/6*2/3